OpenStudy (natalie1234):

PLEASE CHECK IF MY ANSWER IS RIGHT PLEASE Find the number halfway between 3/4 and 3/5 is it 27/40??

OpenStudy (anonymous):

yes, it's correct

OpenStudy (natalie1234):

ok thanks

OpenStudy (anonymous):

are you good with math? i reaLLY need help

OpenStudy (anonymous):

@jamesx3rd yes, i am. :)

OpenStudy (anonymous):

Umm, what is the solution of the following system, x-y=11 -x+y=-11

OpenStudy (anonymous):

is x=y+11 an option? @jamesx3rd

OpenStudy (anonymous):

A.-3, -4 B.No solutions C.infinitely many solutions D.3,4 ;)

OpenStudy (anonymous):

I'm leaning on C...

OpenStudy (anonymous):

thanks